How to Sketch a bode plot for the transfer function given below indicating approximate amplitudes, slopes (in dB/decade) and frequency values (in Hertz).

(s? +4) (52 +220 + 4000)
Answer #1

criven, a(s) = (s² + 4) S2+ 2205 + 4000 = (stj 2) (s-j2) (5 + 20% (5 + 200) = 4(! + 2)(- ) - 20/1 + 2 ) 200(1+2 660) S = jw tInitial slope: Kp a lim (jw) = 0.001 Wyo Tocjus lu- ol = 20 log (0.001) = -60 dB dB At w=2 , (1000g 2 = 3.010) slope = +40 dB